OpenStudy (anonymous):

2. Assume approximately 338 million, or about 30%, of a country’s population is younger than 16. What is the approximate population of the entire country?

OpenStudy (anonymous):

30%(x) = 338 million. solve for x.

hero (hero):

is/of = %/100 338/x = 30/100 33800 = 30x 33800/30 = x 1126.67 = x In other words 1126.67 million is the original pop